OpenStudy (awesomedude2078):

@danieldjpon3 Erving Spencer makes a salary of $24,000 per year and he is paid biweekly. If his employer puts 6.5% into his pension, how much goes into the pension every paycheck?

OpenStudy (danieldjpon3):

Salary of $24,000 per year i.e. for 52 weeks = 461.54 for 1 week = 461.54 × 2 = 923.08 for 2 weeks pension = 923.08 × 6.5/100 = $ 60 in pay check

OpenStudy (danieldjpon3):

I hoe that helps

OpenStudy (danieldjpon3):

hope*